Length of rectangle is 12 inches and its perimeter is6 times its length what isits width
Answer by math_helper(2461)


P = 2L + 2W
You are given L=12in and P=6L:
6L = 2L + 2W
4L = 2W
4*12 = 2W
W = 24in

Check (calculate Perimeter both ways):
P = 2L + 2W = 2*12 + 2*24 = 24 + 48 = 72
P = 6L = 6*12 = 72 (ok)